A firm produces a product that has the production cost function C(X)=330x+15,400 and the revenue function R(X)=440x. No more than 155 units can be sold. Find and analyze the break even quantity then find the Profit function.
Answer by Boreal(15235)

Break-even is when the two functions are equal
330x+15400=440x
110x=15400
x=140 units. At that point, the cost is $46200+$15400=$61600. The revenue is 140*440=$61600
Profit function is 440x-330x-15400 for x<156
For x=150, the profit is $66,000-$49,500-15400=$1100